# EF-Calib: Spatiotemporal Calibration of Event- and Frame-Based Cameras Using Continuous-Time Trajectories

### v1

@@ IV-B2 Feature Extraction
Furthermore, additional geometric constraints are needed to eliminate the remaining false positive regions. Specifically, the two candidate regions constituting the same ellipse should demonstrate similar PCA magnitudes; regions with a notable discrepancy in PCA magnitude fail to meet this criterion. Additionally, the contributions of the two candidate regions to the circumference of the ellipse should be close. In other words, the angular range $\theta_{r}$ of the two candidate regions with respect to the center of the ellipse should be close to 180.
Regions that successfully meet these geometric constraints are recognized as accurately representing the elliptical features within the calibration pattern. Consequently, depending on the distribution of these elliptical features, their relative positions on the calibration pattern can be decoded to correspond with the results identified in the frame-based camera.
@@ IV-B3 Feature Refinement
Neglecting the timestamps of events during elliptical fitting inevitably introduces errors, thereby affecting the camera calibration precision. However, by regarding the timestamps $t$ as a third dimension alongside the pixel coordinates $[x,y]^{T}$, events can be conceptualized as points within a three-dimensional space. Each “accumulation frames”, corresponding to a brief time interval, allows for the assumption of solely translational movement with speed $[v_{x},v_{y}]^{T}$, parallel to the pixel plane, for each elliptical feature at any given moment within this interval, as demonstrated in Fig. 4. Consequently, the moving ellipse model $\mathcal{F}$ is described by the following representation:
$\mathcal{F}:ax(t)^{2}+\beta x(t)y(t)+\gamma y(t)^{2}+\delta x(t)+\epsilon y(t)+\zeta=\text{0}$ (13)
@@
$\begin{cases}\mathcal{F}:\mathbf{P}(t)^{T}\mathbf{Q}\mathbf{P}(t)=\text{0}\\ \mathbf{P}(t)=\mathbf{P}(t_{0})-\mathbf{V}\cdot(t-t_{0})\\ \mathbf{Q}=\begin{pmatrix}\alpha&\beta/\text{2}&\delta/\text{2}\\ \beta/\text{2}&\gamma&\epsilon/\text{2}\\ \delta/\text{2}&\epsilon/\text{2}&\zeta\end{pmatrix}\end{cases}$ (14)
where $\mathbf{P}(t)=[x(t),y(t),\text{1}]^{T}$, $\mathbf{V}=[v_{x},v_{y}]^{T}$, $t\in[t_{0},t_{0}+\text{2}\delta_{t}]$, and $t_{0}$ represents the starting time of current “accumulation frame”. Here the cost function for model optimization is defined as
@@
By substituting the events into the aforementioned model, we utilize the Levenberg-Marquardt algorithm to solve the model, thereby refining the elliptical features.
@@ IV-C Trajectory Optimization
The refinement process converts the previously discrete elliptical features into densely populated patterns within the corresponding time period of associated events. This densely populated feature facilitates the optimization of event camera poses within a continuous-time trajectory representation. Nevertheless, maintaining continuous visibility of the entire calibration plate during the calibration process poses a challenge. Incomplete calibration patterns in a continuous event stream often result in unsuccessful recognition. To mitigate this issue, the continuous-time trajectory is divided into multiple segments based on the output from the recognizer. Consequently, a piece-wise B-spline-based optimizer for event camera pose trajectories is proposed.
@@
where $\xi_{m}^{i}$ is the control point of $\mathcal{L}_{m}$, $N_{m}$ is the number of control points, and $K_{e}$ and $D_{e}$ are the intrinsic and distortion parameters of the event camera. The corresponding visual residual for the $i$th feature based on reprojection error is defined as:
$\mathbf{r}_{e}(\mathcal{X}_{e})=\sum_{k\in\mathcal{K}}\pi_{e}(\mathbf{R}_{w}^{e}(t_{i}+k\cdot\delta_{t_{i}})P_{i}^{w}+\mathbf{t}_{w}^{e}(t_{i}+k\cdot\delta_{t}))$ (18)
$-(\begin{bmatrix}u_{i}^{e}\\ v_{i}^{e}\end{bmatrix}+k\cdot\delta_{t}\mathbf{V}_{i})$
where $\mathbf{R}^{e}_{w}(\cdot)$ and $\mathbf{t}^{e}_{w}(\cdot)$ are derived from the B-spline trajectory using equations (6) and (7), respectively. Since the feature refinement yields a continuous moving ellipse model, the residuals can be constructed by sampling the model at any time. Here, $\mathcal{K}$ denotes the partition of $\delta_{t_{i}}$, which defines the sampling interval of the feature, as shown in Fig. 6. The function $\pi_{e}(\cdot)$ projects the spatial point $P_{i}^{w}$ onto the “accumulation frame”.
@@
where $\rho(\cdot)$ is the Huber loss function.
IV-D Spatialtemporal Calibration
The final step involves jointly optimizing the event camera and the frame-based camera, utilizing the previously optimized trajectories, to determine the extrinsic parameters and time offset. The corresponding state vector $\mathcal{X}_{f}$ is:
@@ IV-D Spatialtemporal Calibration
$\mathbf{r}_{f}(\mathcal{X}_{f})=\pi_{f}\left(\mathbf{R}^{f}_{e}\left(\mathbf{R}^{e}_{w}(t_{i}+t_{d})P^{w}_{i}+\mathbf{t}^{e}_{w}(t_{i}+t_{d})\right)+\mathbf{t}^{f}_{e}\right)-\begin{bmatrix}u^{f}_{i}\\[6.0pt] v^{f}_{i}\end{bmatrix}$ (21)
where $\pi_{e}(\cdot)$ projects the spatial points onto the image plane of the frame-based camera.
@@
$J_{t_{d}}=\frac{\partial\mathbf{r}_{f}}{\partial P_{i}^{f}}\frac{\partial P_{i}^{f}}{\partial t_{d}}=\frac{\partial\mathbf{r}_{f}}{\partial P_{i}^{e}}(\frac{\partial P_{i}^{f}}{\partial\mathbf{R}_{w}^{e}}\frac{\partial\mathbf{R}_{w}^{e}}{\partial t_{d}}+\frac{\partial P_{i}^{f}}{\partial\mathbf{t}_{w}^{e}}\frac{\partial\mathbf{t}_{w}^{e}}{\partial t_{d}})$ (22)
where $P_{i}^{f}=\mathbf{R}^{f}_{e}\left(\mathbf{R}^{e}_{w}(t_{i}+t_{d})P^{w}_{i}+\mathbf{t}^{e}_{w}(t_{i}+t_{d})\right)+\mathbf{t}^{f}_{e}$.
@@
$\frac{\partial P_{i}^{f}}{\partial\mathbf{R}_{w}^{e}}\frac{\partial\mathbf{R}_{w}^{e}}{\partial t_{d}}=\mathbf{R}^{f}_{e}\dot{\mathbf{R}}^{e}_{w}(t_{i}+t_{d})P^{w}_{i}+\mathbf{R}^{f}_{e}\mathbf{v}^{e}_{w}(t_{i}+t_{d})$ (23)
@@
Notably, the optimization problems in (15), (19), and (24) are solved using Google Ceres [^1] https://ceres-solver.org/.
@@ V Experiments
In this section, several experiments are conducted to evaluate the performance of EF-Calib, encompassing intrinsic calibration test, extrinsic calibration test, and time offset calibration test. Additionally, an ablation study is conducted to evaluate the contribution of several key modules within EF-Calib.
@@ V-A System Setup
A real-world stereo vision system was designed as Fig. 1(b) shows. It contains an event camera and a frame-based camera. The two cameras are integrated by a slide, on which the baseline and viewing angle can be arbitrarily changed to test the calibration performance of EF-Calib in different situations comprehensively. The event camera utilized in this letter is the Inivation DAVIS 346, featuring a resolution of 346$\times$260 and a maximum temporal resolution of 1 µs. Additionally, this type of camera can also generate regular frames at a frequency of 30 Hz under standard illumination conditions. This configuration can be readily employed to compare EF-Calib with a high-quality, frame-based calibration pipeline, such as the OpenCV calibration toolbox [29]. The frame-based camera employed is the HikVision MV-CE013-80UM industrial camera with a global shutter and a resolution of 1280$\times$1024 pixels. Note that no hardware synchronization was utilized in the stereo vision system. This deliberate choice was made to provide a more rigorous evaluation of the calibration capability of EF-Calib under real-world conditions.
V-B Calibration Experiments
The calibration performance of a stereo vision system is usually affected by the camera baseline and viewing angle. To fully evaluate the calibration performance of EF-Calib, we conducted calibration experiments in three settings and analyzed the corresponding calibration results separately. In the first setting (Trial 1), the cameras are configured for a regular baseline. In the second setting (Trial 2), the cameras are configured for a wide baseline. In the third setup (Trial 3), the cameras are configured as a narrow baseline. For each trial, the cameras are adjusted to obtain a reasonable viewing angle, ensuring sufficient overlap of the camera field of view. For each trial, images with event data are recorded simultaneously for sufficient time (about 40 s) to achieve converged calibration results.
V-B1 Intrinsic Calibration Test
We utilized only the event stream data from the above three trials and completed the intrinsic calibration of the event camera respectively. Previously, we completed the intrinsic calibration using OpenCV toolbox [29] with the frame provided by DAVIS 346 and considered this calibration result as the ground truth. In addition, EF-Calib is compared with two state-of-the-art event camera intrinsic calibration methods [17, 20]. Note that the calibration patterns used by the compared methods are the ones originally employed by them: [17] utilizes a checkerboard pattern, while [20] employs an asymmetric circular pattern.
Table I shows the intrinsic calibration results of each method. It can be seen that the intrinsic parameter obtained by our method is closest to the ground truth (GT) and the results corresponding to three trials are very stable. In addition, Fig. 7 illustrates the plot of the intrinsic parameters over time. It can be noticed that EF-Calib can get converged results in less than 20 s, demonstrating the ease of use of our method.
@@ V-B2 Extrinsic Calibration Test
To evaluate the extrinsic calibration performance of the EF-Calib, we calculated the errors corresponding to rotation and translation separately for each trial, i.e.
@@
where $N$ is the frame number and $\boldsymbol{\theta}(\cdot)$ represents the Euler angle corresponding to the rotation matrix. Similarly to the intrinsic calibration test, we also acquired 30 pairs of images containing a checkerboard calibration board from the two cameras in different poses simultaneously. These frames were calibrated using the OpenCV toolbox to obtain the corresponding extrinsic parameters for both cameras. The computed EF-Calib calibration errors were compared with the corresponding errors of the frame-based extrinsic parameter calibration. From Table  II, it can be seen that EF-Calib can achieve the same level of error as the frame-based extrinsic calibration, verifying its effectiveness in extrinsic calibration.
@@
Trial 1 Frame-based 0.3499 0.0918 30
EF-Calib 0.5336 0.1984 250
Trial 2 Frame-based 0.4199 0.1854 30
EF-Calib 0.6572 0.3127 207
Trial 3 Frame-based 0.2541 0.0954 30
EF-Calib 0.3638 0.2911 184
@@ V-B3 Time Offset Calibration Test
Time offset estimation is crucial for multi-camera systems that lack hardware triggering. In this test, we evaluate EF-Calib’s capability to calibrate time offsets by manually adjusting the timestamp of each image frame. Specifically, we compare the differences between the time offset calibration results obtained with modified timestamps and the original results. This is achieved by uniformly delaying or advancing the timestamps by a fixed time interval $\Delta t_{d}$. Specifically, the timestamps were modified by $\pm$2.5 ms and $\pm$5 ms for each trial, and the difference between the modified time delay and the original time delay was calculated and compared to the delta. Fig. 8 illustrates the experimental results of the time offset calibration test, and EF-Calib can accurately calibrate the time offset at different scenarios.
Figure 8: Results of the time offset calibration test.
TABLE III: Ablation Study of EF-Calib
Piece-wise trajectory Temporal calibration Feature refinement Trial 1 Trial 2 Trial 3

### v2

@@ IV-B2 Feature Extraction
Furthermore, additional geometric constraints are needed to eliminate the remaining false positive candidate regions. Specifically, the PCA magnitudes of two connected components with opposite polarity, which together form the same ellipse, are calculated separately. For a true ellipse feature, these two semicircular arcs should exhibit similar PCA magnitudes; therefore, candidates with significant differences in PCA magnitudes are considered false positives. Additionally, the contribution of both connected components to the ellipse’s circumference should be nearly equal. This means that the angular range $\theta_{r}$ of the two candidate regions, relative to the center of the ellipse, should be close to 180.
Regions that successfully meet these geometric constraints are considered to accurately represent the elliptical features in the calibration pattern. The next step involves decoding the relative positions of the elliptical features using their spatial distribution to align with the frame-based camera’s recognized results. The calibration pattern’s default orientation is set so that the upper-left feature has black regions in the upper-left and lower-right directions. By identifying the elliptical feature in the upper-left corner of the current “accumulation frames”, the pattern’s orientation can be determined from the event distribution.
@@ IV-B3 Feature Refinement
Neglecting the timestamps of events during elliptical fitting inevitably introduces errors, thereby affecting the precision of the calibration results. However, by regarding the timestamps $t$ as a third dimension alongside the pixel coordinates $[x,y]^{T}$, events can be conceptualized as points within a three-dimensional space. Each “accumulation frames”, corresponding to a brief time interval, allows for the assumption of solely translational movement with speed $[v_{x},v_{y}]^{T}$, parallel to the pixel plane, for each elliptical feature at any given moment within this interval, as demonstrated in Fig. 4. Consequently, the moving ellipse model $\mathcal{F}$ is described by the following representation:
$\mathcal{F}:\alpha x(t)^{2}+\beta x(t)y(t)+\gamma y(t)^{2}+\eta x(t)+\epsilon y(t)+\zeta=\text{0}$ (9)
@@
$\begin{cases}\mathcal{F}:\mathbf{P}(t)^{T}\mathbf{Q}\mathbf{P}(t)=\text{0}\\ \mathbf{P}(t)=\mathbf{P}(t_{0})-\mathbf{V}\cdot(t-t_{0})\\ \mathbf{Q}=\begin{pmatrix}\alpha&\beta/\text{2}&\eta/\text{2}\\ \beta/\text{2}&\gamma&\epsilon/\text{2}\\ \eta/\text{2}&\epsilon/\text{2}&\zeta\end{pmatrix}\end{cases}$ (10)
where $\mathbf{P}(t)=[x(t),y(t),\text{1}]^{T}$, $\mathbf{V}=[v_{x},v_{y}]^{T}$, $t\in[t_{0},t_{0}+\text{2}\delta_{t}]$, and $t_{0}$ represents the starting time of current “accumulation frame”. Notice that the moving ellipse model is optimized separately for each elliptic feature, and the cost function for model optimization is defined as
@@
An initial moving ellipse model is created by first calculating the average timestamps of the events and fitting the elliptical model using only their spatial coordinates. Then, by substituting the events into this model, we refine the elliptical features through optimization using the Levenberg-Marquardt algorithm.
@@ IV-C Trajectory Optimization
After obtaining the refined features, the calibration parameters are optimized to yield the final results. Given the additional degrees of freedom introduced by the continuous-time trajectory, simultaneously optimizing all parameters risks the system becoming trapped in local minima. To address this, a two-stage optimization strategy is employed. Initially, the intrinsic parameters and motion trajectory of the event camera are optimized. These optimized parameters then serve as prior information to refine the remaining variables. The validity and superiority of this approach are demonstrated through the ablation study in Sec. V.
The feature refinement process turns discrete elliptical features into densely populated patterns within the relevant time period, facilitating the optimization of event camera poses in a continuous-time trajectory. However, continuous visibility of the calibration pattern during the process is difficult to maintain, often resulting in recognition failures due to incomplete patterns in the event stream. To counter this, the trajectory is segmented based on recognizer output, and a piecewise B-spline-based optimizer for event camera pose trajectories is introduced.
@@
where $\xi_{m}^{i}$ is the control point of $\mathcal{L}_{m}$, $N_{m}$ is the number of control points, and $K_{e}$ and $D_{e}$ are the intrinsic and distortion parameters of the event camera. The corresponding visual residual for the $j$th feature based on reprojection error is defined as:
$\mathbf{r}_{e}(\mathcal{X}_{e})=\sum_{k\in\mathcal{K}}\pi_{e}(\mathbf{R}_{w}^{e}(t_{i}+k\cdot\delta_{t_{i}})P_{j}^{w}+\mathbf{t}_{w}^{e}(t_{i}+k\cdot\delta_{t_{i}}))$ (14)
$-(\begin{bmatrix}u_{j}^{e}\\ v_{j}^{e}\end{bmatrix}+k\cdot\delta_{t_{i}}\mathbf{V}_{i})$
where $\mathbf{R}^{e}_{w}(\cdot)$ and $\mathbf{t}^{e}_{w}(\cdot)$ are derived from the B-spline trajectory using (5) and (6), respectively, and $P_{j}^{w}$ is the 3D spatial point of the $j$th feature in the world coordinate system, obtained from the geometric relationship of the calibration pattern. Since the feature refinement yields a continuous moving ellipse model, the residuals can be constructed by sampling the model at any time. Here, $\mathcal{K}$ denotes the partition of $\delta_{t_{i}}$, which defines the sampling interval of the feature, as shown in Fig. 6. The function $\pi_{e}(\cdot)$ projects the spatial point $P_{j}^{w}$ onto the “accumulation frame”.
@@
where $\rho(\cdot)$ is the Huber loss function, employed to reduce the impact of outliers. In the optimization problem, only the residuals corresponding to features whose timestamps lie within these piecewise trajectories are incorporated.
IV-D Spatiotemporal Calibration
The final step involves optimizing the intrinsic and extrinsic parameters of the frame-based camera, as well as the time offset between the two cameras, using the previously optimized trajectories. The corresponding state vector $\mathcal{X}_{f}$ is:
@@ IV-D Spatiotemporal Calibration
$\mathbf{r}_{f}(\mathcal{X}_{f})=\pi_{f}\left(\mathbf{R}^{f}_{e}\left(\mathbf{R}^{e}_{w}(t_{i}+t_{d})P^{w}_{j}+\mathbf{t}^{e}_{w}(t_{i}+t_{d})\right)+\mathbf{t}^{f}_{e}\right)-\begin{bmatrix}u^{f}_{j}\\[6.0pt] v^{f}_{j}\end{bmatrix}$ (17)
where $\mathbf{R}^{e}_{w}(\cdot)\in\mathbb{SO}$3 and $\mathbf{t}^{e}_{w}(\cdot)\in\mathbb{SO}$3 refer to the rotation and translation of the event camera pose represented by the B-spline, and $\pi_{f}(\cdot)$ projects the spatial points onto the image plane of the frame-based camera.
@@
$J_{t_{d}}=\frac{\partial\mathbf{r}_{f}}{\partial P_{j}^{f}}\left(\frac{\partial P_{j}^{f}}{\partial t_{d}}\right)=\frac{\partial\mathbf{r}_{f}}{\partial P_{j}^{e}}\left(\frac{\partial P_{j}^{f}}{\partial\mathbf{R}_{w}^{e}}\frac{\partial\mathbf{R}_{w}^{e}}{\partial t_{d}}+\frac{\partial P_{j}^{f}}{\partial\mathbf{t}_{w}^{e}}\frac{\partial\mathbf{t}_{w}^{e}}{\partial t_{d}}\right)$ (18)
where $P_{j}^{f}=\mathbf{R}^{f}_{e}\left(\mathbf{R}^{e}_{w}(t_{i}+t_{d})P^{w}_{i}+\mathbf{t}^{e}_{w}(t_{i}+t_{d})\right)+\mathbf{t}^{f}_{e}$.
@@
$\frac{\partial P_{j}^{f}}{\partial\mathbf{R}_{w}^{e}}\frac{\partial\mathbf{R}_{w}^{e}}{\partial t_{d}}=\mathbf{R}^{f}_{e}\dot{\mathbf{R}}^{e}_{w}(t_{i}+t_{d})P^{w}_{j}+\mathbf{R}^{f}_{e}\mathbf{v}^{e}_{w}(t_{i}+t_{d})$ (19)
@@
In the optimization process, an initial value must be provided. Here, the initial value for the time offset $t_{d}$ is set to 0. We then iterate through the timestamps of each frame from the frame-based camera and each “accumulation frame” from the event camera, identifying the pair with the closest timestamps. This pair is used to calculate for the corresponding extrinsic parameters, which are then used as the initial $\mathbf{T}^{f}_{e}$. Notably, the optimization problems in (11), (15), and (20) are solved using Google Ceres [31].
@@ V Experiments
In this section, several experiments are conducted to evaluate the performance of EF-Calib, encompassing intrinsic calibration test, extrinsic calibration test, and time offset calibration test. Additionally, two ablation studies are conducted to evaluate the contribution of several key modules and the optimization strategy within EF-Calib.
@@ V-A System Setup
A real-world stereo vision system was designed as Fig. 1(b) shows. It contains an event camera and a frame-based camera. The two cameras are integrated by a slide, on which the baseline and viewing angle can be arbitrarily adjusted to test the calibration performance of EF-Calib in different situations comprehensively. The event camera utilized in this letter is the Inivation DAVIS 346, featuring a resolution of 346$\times$260 and a maximum temporal resolution of 1 µs. Additionally, this type of camera can also generate regular frames at a frequency of 30 Hz under standard illumination conditions. This configuration can be readily employed to compare EF-Calib with a high-quality, frame-based calibration pipeline, such as the OpenCV calibration toolbox [32]. The frame-based camera employed is the HikVision MV-CE013-80UM industrial camera with a global shutter and a resolution of 1280$\times$1024 pixels. Note that no hardware synchronization was utilized in the stereo vision system. This deliberate choice was made to provide a more rigorous evaluation of the calibration capability of EF-Calib under real-world conditions.
@@
V-B Calibration Experiments
The calibration performance of a stereo vision system is usually affected by the camera baseline and viewing angle. To fully evaluate the calibration performance of EF-Calib, we conducted calibration experiments in three settings and analyzed the corresponding calibration results separately. In the first setting (Trial 1), the cameras are configured for a regular baseline. In the second setting (Trial 2), the cameras are configured for a wide baseline. In the third setup (Trial 3), the cameras are configured as a narrow baseline. For each trial, the cameras are adjusted to obtain a reasonable viewing angle, ensuring sufficient overlap of the camera FOV. For each trial, images with event data are recorded simultaneously for sufficient time ($\sim$40 s) to achieve converged calibration results.
V-B1 Intrinsic Calibration Test
We utilized the event stream data from each of the three trials mentioned above to independently complete the intrinsic calibration of the event camera. Previously, we completed the intrinsic calibration using OpenCV toolbox [32] with the frame provided by DAVIS 346 and considered this calibration result as the ground truth. We obtained the intrinsic parameters and reprojection errors (RPE) of the event camera using EF-Calib and two state-of-the-art intrinsic calibration methods [18, 21], and compared them with the ground truth. Note that the calibration patterns used by the compared methods are the ones originally employed by them: [18] utilizes a checkerboard pattern, while [21] employs an asymmetric circular pattern.
Table I shows the intrinsic calibration results for each method. As shown, the intrinsic parameters derived from our method are the closest to the ground truth (GT), with consistently low RPE across all trials, and the results are highly stable across the three trials. In addition, Fig. 7 illustrates the plot of the intrinsic parameters over time. It can be noticed that EF-Calib can get converged results in less than 20 s, demonstrating the ease of use of our method.
@@ V-B2 Extrinsic Calibration Test
To evaluate the extrinsic calibration performance of the EF-Calib, we calculated the errors corresponding to rotation and translation separately for each trial, i.e.,
@@
where $N$ is the frame number and $\boldsymbol{\theta}(\cdot)$ represents the absolute angle corresponding to the rotation matrix. Similar to the intrinsic calibration test, we also acquired 30 pairs of images (each is captured statically, comprising one image from the frame-based camera and one image from the DAVIS 346) featuring a checkerboard calibration board, captured simultaneously from different viewing angles by the two cameras. These image pairs were then calibrated using the OpenCV toolbox to obtain the corresponding extrinsic parameters for both cameras. We refer to this method as “frame-based”. According to (21), we can calculate the rotation and translation errors corresponding to the “frame-based” method and use them as a reference to evaluate the extrinsic calibration accuracy of EF-Calib. From Table II, it can be seen that EF-Calib can achieve the same level of error as the frame-based extrinsic calibration, verifying its effectiveness in extrinsic calibration.
@@
Trial 1 Frame-based 0.350 0.092 30
EF-Calib 0.534 0.198 250
Trial 2 Frame-based 0.420 0.185 30
EF-Calib 0.657 0.313 207
Trial 3 Frame-based 0.254 0.095 30
EF-Calib 0.364 0.291 184
Figure 8: Results of Time Offset Calibration Test. (a) Without hardware synchronization. (b) With hardware synchronization.
@@ V-B3 Time Offset Calibration Test
Time offset estimation is crucial for multi-camera systems without hardware triggers. In this test, we evaluated EF-Calib’s ability to calibrate the time offset. First, we compared the results with the original calibration by manually adjusting the timestamps. Specifically, the timestamps were modified by $\pm$2.5 ms and $\pm$5 ms for each trial, and the differences between the adjusted and original time delays were calculated and compared to the $\Delta t_{d}$. As shown in Fig. 8(a), EF-Calib accurately calibrates the time offset across different trials. Additionally, we synchronized the cameras’ timestamps via hardware and then calibrated the time offset under the same settings to see if $t_{d}$ approaches 0 ms. Fig. 8(b) shows that EF-Calib successfully calibrates the time offset to near 0 ms after hardware synchronization.
V-C Ablation Studies
TABLE III: Ablation Study of the Key Modules in EF-Calib
Piece-wise trajectory Temporal calib Feature refine Trial 1 Trial 2 Trial 3